Platinum has a density of 21.4 g/cubic centimeters. 900 pounds are stolen. How many 1 liter bars of platinum could they take?
1
Expert's answer
2020-08-25T12:18:40-0400

"V=\frac{m}{\rho}"

density of 21.4 g/cubic centimeters

m=900 pounds=408233 g

"V=\frac{m}{\rho}=\frac{408233}{21.4}=19076.31" cubic centimeters

1 cubic centimeter equals 0.001 liter

V=19,07631 liter=19 liter